OpenStudy (anonymous):

John and alan have a collection of x baseball cards. John has x/4 cards. What fraction of the cards does Alan have

OpenStudy (anonymous):

\[ \text{Alan's cards} = \text{All cards} - \text{John's cards} = x - \frac{x}{4} = \frac{4x}{4} - \frac{x}{4} = \frac{3x}{4}\]